OpenStudy (anonymous):

a and b are positive integers and: f(x) = x^n(a - bx)^n 1) What are the max and min values of function on interval [0, a/b] 2) Use the max value to give upper value for integral from 0 to a/b of f(x) 3) Explain why it is possible to find an n such that the integral in part 2 is less than 1

OpenStudy (anonymous):

1) Since f_n(x) = (ax - bx^2)^n, f '_n(x) = n(ax - bx^2)^(n-1) * (a - 2bx). Setting this equal to 0 yields x = 0, a/b, a/(2b). Testing these values (noting that x = 0, a/b are the endpoints of the interval as well): f_n(0) = f_n(a/b) = 0 <----Minimum f_n(a/(2b)) = (a^2/(4b))^n <----Maximum --------------- 2) ∫(x = 0 to a/b) f_n(x) dx < ∫(x = 0 to a/b) (a^2/(4b))^n dx = (a^2/(4b))^n * (a/b). 3) If we are assuming that a^2 < 4b (or even just a < b), then a^2/(4b) < 1 and hence (a^2/(4b))^n tends to 0 as n goes to infinity. Hence, (a^2/(4b))^n * (a/b) < 1 for sufficiently large n.